# Bézier Splatting for Fast and Differentiable Vector Graphics

### v2

@@ Abstract
Differentiable vector graphics (VGs) are widely used in image vectorization and vector synthesis, while existing representations are costly to optimize and struggle to achieve high-quality rendering results for high-resolution images. This work introduces a new differentiable VG representation, dubbed Bézier splatting, that enables fast yet high-fidelity VG rasterization. Bézier splatting samples 2D Gaussians along Bézier curves, which naturally provide positional gradients at object boundaries. Thanks to the efficient splatting-based differentiable rasterizer, Bézier splatting achieves over 20× and 150× faster per forward and backward rasterization step for open curves compared to DiffVG. Additionally, we introduce an adaptive pruning and densification strategy that dynamically adjusts the spatial distribution of curves to escape local minima, further improving VG quality. Experimental results show that Bézier splatting significantly outperforms existing methods with better visual fidelity and 10× faster optimization speed.
Figure 1: This work proposes Bézier splatting, a new differentiable vector graphic (VG) representation that enables fast and high-fidelity vectorization of high-resolution images, achieving the outstanding rendering performance, 20$\times$ forward computation speedup, 150$\times$ backward computation speedup, and over 10$\times$ total training speedup (tested on a NVIDIA A100 GPU for open curve) in comparison with the state-of-the-art differentiable VG rasterization methods. DiffVG [13] and Bézier splatting support both open and closed curves, while LIVE [16] supports closed curves only. Zoom in for a better view of details.
@@ 1 Introduction
This work presents Bézier splatting, a new differentiable VG representation that optimizes Bézier curves through Gaussian splatting-based rasterization. We sample 2D Gaussian points along Bézier curves and their interior regions, then leverage the Gaussian splatting framework [10] for efficient curve rasterization. Unlike DiffVG [13] which requires computationally intensive boundary sampling and gradient computation, 2D Gaussians inherently provide direct positional gradients at object boundaries through its differentiable Gaussian formulation, enabling over 150$\times$ faster backward computation over DiffVG for open curves. We further introduce an adaptive pruning and densification strategy that adaptively removes redundant curves while adding new ones to necessary regions during optimization. It helps the optimization process escape the local minima of current spatial distributions of curves, formulating a “global receptive field” for further improving the VG optimization process. As shown in Fig. 1, Bézier splatting outperforms existing differentiable VG rasterizers for both open and closed curves. The open curve training takes only $2.9$ minutes, representing an over 10$\times$ training acceleration compared to DiffVG [13] (43 minutes). Bézier splatting preserves better image details and captures fine-grained texture information. For instance, in Fig. 1(a), Bézier splatting presents richer facial details. While LIVE [3] enhances details to some extent, its layer-wise optimization strategy results in artifacts like non-smooth shape boundaries on the sky as well as significantly more computing cost, whereas Bézier splatting enables 88$\times$ training acceleration. We summarize the contribution of this work as follows.
  • We propose a novel differentiable vector graphic representation, Bézier splatting, which achieves 20$\times$ faster forward and 150$\times$ faster backward computation, while producing higher-fidelity VGs for high-resolution images.
@@
  • Extensive experiments demonstrate that Bézier splatting outperforms existing differentiable VG representations in efficiency and visual quality.
@@ 2.1 Vectorization and Rasterization
Image vectorization and vector graphic (VG) rasterization are important research topics in computer graphics and computer vision. DiffVG [13] introduces the first differentiable VG rasterization framework, laying a foundation for optimizing and generating vectorized representations through gradient-based methods. It expands the applicability of VGs to a broader range of tasks including image vectorization, text-to-VG generation, and painterly rendering. Based on DiffVG [13], LIVE [16] further proposes a layer-wise path initialization strategy, vectorizing raster images into compact and semantically consistent VG representations while preserving image topology. Du et al. [3] use linear gradients to fill the regional colors, and Chen et al. [2] propose a specific implicit neural representation to model regional color distributions, enhancing the color representation within closed Bézier curves. However, these DiffVG-based VG rasterization approaches [13, 16, 3, 2] suffer from slow optimization, often requiring several hours to process a 2K-resolution image with 1,024 curves.
@@ 3.1 Overall
To overcome this inefficiency, this work proposes a novel VG representation, Bézier splatting, which is inspired by the high computational efficiency and expressive fitting capacity of Gaussian splatting [32] for rasterization. Bézier splatting samples points along Bézier curves and represents them as 2D Gaussians, enjoying the following advantages:
  • This design significantly accelerates the forward and backward pass of Bézier curve rasterization, achieving over 10× training speedup without specialized optimization techniques;
@@ 3.2 Primitives of Bézier Splatting
Each Bézier curve $\mathcal{B}_{i}(t)$ is associated with an RGB color parameter $c_{i}\in\mathbb{R}^{3}$, and an opacity parameter ø$\o_{i}\in[0,1]$ that defines the transparency of the curve.
@@ 3.3 Sampling Gaussians on Bézier Curves
To address this limitation, we propose a new structure named paired Bézier curve structure that facilitates an efficient closed area sampling between two Bézier curves of a closed curve. Specifically, the two Bézier curves $\mathcal{B}_{1}(t)$ and $\mathcal{B}_{R+1}(t)$ share the same start and end points, and $R$ intermediate Bézier curves are generated by interpolating the control points between the two Bézier curves $P^{(0)}_{j}$ and $P^{(R+1)}_{j}$as:
$P^{(k)}_{j}=(1-t_{k})P^{(0)}_{j}+t_{k}P^{(R+1)}_{j},\;k=1,\dots,R,$ (4)
@@
Let $\mathbf{X}_{r,k}$ represents the 2D position of the $k$-th sampled point on the $r$-th interpolated curve, the x-direction scale $\sigma_{x}(r,k)$ of a 2D Gaussian is calculated by the pairwise distance between consecutive points along the curve, and y-direction scale $\sigma_{y}(r,k)$ is calculated by the pairwise distance between corresponding points on adjacent curves:
$\sigma_{x}(r,k)$ $=\|\mathbf{X}_{r,k+1}-\mathbf{X}_{r,k}\|_{2}/\rho,$ (7)
$\sigma_{y}(r,k)$ $=\|\mathbf{X}_{r+1,k}-\mathbf{X}_{r,k}\|_{2}/\rho.$
$\rho$ is a constant that determines the density of 2D Gaussian points. The rotation $\theta$ of each Gaussian point is calculated by its left and right neighbors, as
$\theta_{r,k}=\operatorname{atan2}\left(y_{r,k+1}-y_{r,k-1},x_{r,k+1}-x_{r,k-1}\right)$ (8)
For points on boundaries, the rotation is set to align with the nearest neighbor.
For both closed and open curves, the depth $d$ of a Gaussian is assigned based on the area of curves, ensuring that smaller curves are not occluded by larger ones. This prevents them from being ignored during optimization, as the gradient would not count Gaussians when the accumulated opacity exceeds $1$.
@@ 3.4 Differentiable Rasterization by Splatting
Discussion on efficiency. The rasterization process of Bézier splatting is very efficient, since we directly sample 2D Gaussians from Bézier curves in a differentiable manner, then splatting them to the 2D plane. The rasterization pipeline remains end-to-end differentiable while being highly optimized for parallel computation and large-scale matrix operations. As a result, Bézier splatting is a highly efficient and fully differentiable VG representation, which ensures both fast rendering and gradient-based optimization. It not only preserves the flexibility of VGs but also allows seamless integration into deep learning frameworks, making it well-suited for tasks requiring high-quality and editable vector representations.
@@ 3.5 Optimization
Training objective. Given a raster image $\mathcal{I}\in\mathbb{R}^{H\times W\times 3}$, our goal is to vectorize it into Bézier curves while achieving high-fidelity reconstruction. We first randomly generate a set of Bézier curves to lay on the canvas, then employ the differentiable rasterizer discussed in Sec. 3.4 to render a raster image ^$\hat{\mathcal{I}}\in\mathbb{R}^{H\times W\times 3}$. The Bézier curves can then be optimized through any gradient-based loss functions. In this work, we formulate the optimization objective as minimizing the loss between $\mathcal{I}$ and $\hat{\mathcal{I}}$ while enforcing the curves to be convex. Therefore, we only adopt an $L_{2}$ loss and a Xing loss $L_{\text{Xing}}$ [16], as:
@@
Figure 3: We showcase the image vectorization results by our Bézier splatting, with open or closed curves and different curve counts (indicated at the bottom right of each example). The raster inputs, shown at the top left of each example, are taken from the DIV2K [22], KIOAK [12], Clipart [8], and Danbooregions [29] datasets. Zoom in for a better view of details.
Adaptive curve pruning and densification. The gradients of 2D Gaussians are influenced by local pixels only. Therefore, it is hard for 2D Gaussians to dynamically reallocate to regions requiring finer-grained details. The Gaussians would be trapped in local minima, leading to redundant Gaussians that are optimized as either low opacity or excessively large size, resulting in artifacts in the rendering result.
@@
Figure 4: A qualitative comparison of our method and the existing differentiable VG rasterization methods, including DiffVG [13] and LIVE [16], with different number ($N$) of open and closed curves. The raster input is taken from the DIV2K dataset [22].
To address this issue, this work introduces a new pruning and densification strategy to dynamically adjust the density of Bézier curves throughout the optimization process. For pruning redundant Bézier curves, we apply three criteria to ensure a stable and precise optimization process. First, we remove curves with opacity below a dynamic threshold that gradually decreases as optimization progresses, ensuring that weakly contributing curves are eliminated while preserving essential structures. Second, we remove curves with an area below a predefined threshold, as they contribute minimally to the final representation. Additionally, we eliminate curves where the opacity is significantly lower in the middle than at both ends. This phenomenon occurs when a single curve improperly spans a region that should be represented by multiple curves, causing the optimization to reduce opacity in the middle while maintaining high opacity at the edges. Third, we remove curves that exhibit high color similarity with surrounding curves and have significant overlap, as they provide little additional information and can be pruned without affecting the final representation. By removing such curves, we allow the optimization to introduce more appropriate replacements, leading to a more accurate and visually coherent vectorized representation. For curve densification, we adopt an error-driven curve allocation strategy inspired by LIVE [16]. Specifically, we compute connected error regions, rank them by area, and add new Bézier curves into the highest-error regions. This adaptive redistribution mechanism ensures that curves are allocated to where they would contribute the most to the reconstruction fidelity. The pruning and densification strategy enables a “global receptive field” for redistributing curve density, effectively preserving visual details for high-fidelity rendering while maintaining a compact vector representation.
@@ 4.1 Experimental Setups
Implementation details. We implement Bézier splatting in PyTorch [18] and optimize it by using the Adam optimizer [11] with a StepLR learning rate scheduler. The learning rate is initialized at 0.01 for color, 2e-4 for Bézier curve control points, and 0.1 for opacity. For the pruning and densification strategy, the opacity threshold is set to 0.02, and the overlap threshold based on Axis-Aligned Bounding Boxes (AABB) is set to 0.9. Following the approach in LIVE [16], new curves are initialized in a circular pattern, and the number of added curves matches the number of removed ones to maintain a constant total curve count. Open curves are optimized for 15,000 iterations, and closed curves are optimized for 10,000 iterations. Pruning and densification are applied every 400 steps until 14,000 or 9,200 steps, after which they are halted to stabilize the representation.
Datasets. We comprehensively evaluate our method across different image domains. We use the publicly available DIV2K [22] dataset for evaluating natural images. Due to the high computational cost of baseline method LIVE [16], we uniformly subsample the DIV2K dataset by selecting one out of every four images, resulting in a final evaluation set of 200 images from the original 800-image dataset. Additionally, we evaluate our method on non-photorealistic images, including the clip art images from Clipart1K dataset [8], and cartoon images from Danbooregions dataset [29], to demonstrate its effectiveness for diverse types of images.
Table 1: Forward and backward speed of DiffVG [13] and Bézier splatting (tested on a 2,040$\times$1,344 image, using 2,048 curves).
DiffVG Bézier splatting Speedup
Open Forward 141.3ms 7.2ms 19.6$\times$
Backward 701.3.ms 4.7ms 149.2$\times$
Closed Forward 85.2ms 14.2ms 6.0$\times$
Backward 448.3ms 15.7ms 28.5$\times$
Table 2: Quantitative evaluation and training efficiency of differentiable VG methods on the DIV2K dataset [22] (256 to 2,048 curves).
Method 256 512 1024 2048
$SSIM^{\uparrow}$ $PSNR^{\uparrow}$ $LPIPS^{\downarrow}$ Train $SSIM^{\uparrow}$ $PSNR^{\uparrow}$ $LPIPS^{\downarrow}$ Train $SSIM^{\uparrow}$ $PSNR^{\uparrow}$ $LPIPS^{\downarrow}$ Train $SSIM^{\uparrow}$ $PSNR^{\uparrow}$ $LPIPS^{\downarrow}$ Train
Open DiffVG 0.552 19.83 0.563 18m56s 0.587 21.47 0.537 22m01s 0.616 22.62 0.517 30m38s 0.654 23.89 0.492 46m36s
Ours 0.600 22.17 0.540 3m37s 0.646 23.79 0.498 3m21s 0.699 25.45 0.448 3m17s 0.756 27.22 0.389 3m24s
Closed DiffVG 0.578 20.69 0.548 16m27s 0.601 21.82 0.531 18m35s 0.631 22.95 0.509 25m08s 0.668 24.21 0.482 34m16s
LIVE 0.576 20.09 0.543 2h41m 0.611 21.70 0.521 4h12m 0.648 23.11 0.495 5h6m - - - -
Ours 0.581 20.74 0.547 2m22s 0.611 22.14 0.522 2m45s 0.648 23.55 0.491 3m25s 0.690 24.93 0.454 4m46s
Figure 5: Ablation study on the proposed adaptive pruning and densification strategy and multi-opacity strategy.

### v3

@@ Abstract
Differentiable vector graphics (VGs) are widely used in image vectorization and vector synthesis, while existing representations are costly to optimize and struggle to achieve high-quality rendering results for high-resolution images. This work introduces a new differentiable VG representation, dubbed Bézier Splatting, that enables fast yet high-fidelity VG rasterization. Bézier Splatting samples 2D Gaussians along Bézier curves, which naturally provide positional gradients at object boundaries. Thanks to the efficient splatting-based differentiable rasterizer, Bézier Splatting achieves 30× and 150× faster per forward and backward rasterization step for open curves compared to DiffVG. Additionally, we introduce an adaptive pruning and densification strategy that dynamically adjusts the spatial distribution of curves to escape local minima, further improving VG quality. Furthermore, our new VG representation supports conversion to standard XML-based SVG format, enhancing interoperability with existing VG tools and pipelines. Experimental results show that Bézier Splatting significantly outperforms existing methods with better visual fidelity and significant optimization speedup. The project page is bezier-splatting.github.io.
Figure 1: This work proposes Bézier Splatting, new differentiable vector graphics (VGs) representation that achieve an order-of-magnitude computational speedup (tested on a NVIDIA RTX 4090 GPU) in comparison with the state-of-the-art differentiable VGs rendering method DiffVG [14].
@@ 1 Introduction
This work presents Bézier Splatting, a new differentiable VG representation that optimizes Bézier curves through Gaussian splatting-based rasterization. We sample 2D Gaussian points along Bézier curves and their interior regions, then leverage the Gaussian Splatting framework [11] for efficient curve rasterization. Unlike DiffVG [14] which requires computationally intensive boundary sampling and gradient computation, 2D Gaussians inherently provide direct positional gradients at object boundaries through its differentiable Gaussian formulation, enabling over 150$\times$ faster backward computation over DiffVG for open curves. We further introduce an adaptive pruning and densification strategy that adaptively removes redundant curves while adding new ones to necessary regions during optimization. It helps the optimization process escape the local minima of current spatial distributions of curves, formulating a “global receptive field” for further improving the VG optimization process. As shown in Fig. 1, Bézier Splatting outperforms the state-of-the-art differentiable VG rasterizer DiffVG [14] for both open and closed curves in terms of efficiency and rendering quality. We summarize the contribution of this work as follows.
  • We propose a novel differentiable vector graphic representation, Bézier Splatting, which achieves 30$\times$ faster forward and 150$\times$ faster backward computation, while producing high-quality rendering results.
@@
  • Extensive experiments demonstrate that Bézier Splatting outperforms existing differentiable VG rendering methods in efficiency and visual quality.
@@ 2.1 Vectorization and Rasterization
Image vectorization and vector graphic (VG) rasterization are important research topics in computer graphics and computer vision. DiffVG [14] introduces the first differentiable VG rasterization framework, laying a foundation for optimizing and generating vectorized representations through gradient-based methods. It expands the applicability of VGs to a broader range of tasks including image vectorization, text-to-VG generation, and painterly rendering. Based on DiffVG [14], LIVE [17] and O&R [6] further proposes a layer-wise path initialization strategy, vectorizing raster images into compact and semantically consistent VG representations while preserving image topology. Du et al. [3] use linear gradients to fill the regional colors, and Chen et al. [2] propose a specific implicit neural representation to model regional color distributions, enhancing the color representation within closed Bézier curves. However, these DiffVG-based VG rasterization approaches [14, 17, 3, 2] suffer from slow optimization, often requiring several hours to process a 2K-resolution image with 1,024 curves.
@@ 3.1 Overall
To overcome this inefficiency, this work proposes a novel VG representation, Bézier Splatting, which is inspired by the high computational efficiency and expressive fitting capacity of Gaussian splatting [34] for rasterization. Bézier Splatting samples 2D Gaussian points along Bézier curves and their interior regions, then leverages the Gaussian splatting method for efficient rasterization, enjoying the following advantages:
  • This design significantly accelerates the forward and backward pass of Bézier curve rasterization, achieving an order-of-magnitude speedup without specialized optimization techniques;
@@ 3.2 Primitives of Bézier Splatting
Each Bézier curve $\mathcal{B}_{i}(t)$ is associated with an RGB color parameter $c_{i}\in\mathbb{R}^{3}$, and an opacity parameter $o_{i}\in[0,1]$ that defines the transparency of the curve.
@@ 3.3 Sampling Gaussians on Bézier Curves
For both closed and open curves, the depth $d$ of a Gaussian is assigned based on the area of curves, ensuring smaller curves are not occluded by larger ones. It prevents them from being ignored during optimization, as the gradient would not count Gaussians when the accumulated opacity exceeds 1.
@@
To address this limitation, we propose a new structure named paired Bézier curve structure, which supports two groups of Bézier curves with equal cardinality and arbitrary degrees, forming a closed region. This structure enables efficient and flexible sampling within the enclosed region between the two curve groups. Specifically, the two Bézier curves $\mathcal{B}_{1}(t)$ and $\mathcal{B}_{R+1}(t)$ share the same start and end points, forming the inside curves and boundaries of a closed region. A total of $R$ intermediate Bézier curves are generated by linearly interpolating the corresponding control points $P^{(0)}_{j}$ and $P^{(R+1)}_{j}$ as:
$P^{(k)}_{j}=(1-t_{k})P^{(0)}_{j}+t_{k}P^{(R+1)}_{j},\quad k=1,\dots,R,$ (4)
where $t_{k}\in[0,1]$ are sampled from a normalized cumulative distribution function (CDF). The interpolated control points $\{P^{(k)}_{j}\}$ define intermediate curves that form a dense strip between the two boundaries. Note that this paired structure can naturally extend to cases where each curve is composed of multiple connected Bézier segments, provided that all segments have the same number of control points, similar to the representation used in DiffVG [14].
@@
Let $\mathbf{X}_{r,k}$ denote the 2D position of the $k$-th sampled point on the $r$-th interpolated curve. To construct anisotropic 2D Gaussian primitives aligned with the local geometry, we define the spatial scales as follows. The x-direction of each 2D Gaussian is defined to align with the local tangent direction of the curve (i.e., along the curve), while the y-direction is perpendicular to it (i.e., across adjacent curves).
The scale $\sigma_{x}(r,k)$ is computed from the Euclidean distance between consecutive points along the curve, and $\sigma_{y}(r,k)$ is computed from the distance between corresponding points on adjacent curves:
$\sigma_{x}(r,k)$ $=|\mathbf{X}_{r,k+1}-\mathbf{X}_{r,k}|_{2}/\rho,$ (7)
$\sigma_{y}(r,k)$ $=|\mathbf{X}_{r+1,k}-\mathbf{X}_{r,k}|_{2}/\rho.$
Here, $\rho$ is a global constant that controls the overall density and overlap of the Gaussians. The rotation $\theta_{r,k}$ of each Gaussian is defined by the angle of the local tangent vector, estimated from neighboring points as
$\theta_{r,k}=\operatorname{atan2}\left(y_{r,k+1}-y_{r,k-1},x_{r,k+1}-x_{r,k-1}\right).$ (8)
For boundary points, the rotation is set to align with the nearest available neighbor.
@@ 3.4 Differentiable Rasterization by Splatting
Discussion on efficiency. The rasterization process of Bézier Splatting is very efficient, since we directly sample 2D Gaussians from Bézier curves in a differentiable manner, then splat them to the 2D plane. The rasterization pipeline remains end-to-end differentiable while being highly optimized for parallel computation and large-scale matrix operations. As a result, Bézier Splatting is a highly efficient and fully differentiable VG representation, which ensures both fast rendering and gradient-based optimization. It not only preserves the flexibility of VGs but also allows seamless integration into deep learning frameworks, making it well-suited for tasks requiring high-quality and editable vector representations.
@@ 3.5 Optimization
Training objective. Given a raster image $\mathcal{I}\in\mathbb{R}^{H\times W\times 3}$, the goal of image vectorization is to vectorize the image into Bézier curves while ensuring a high-fidelity reconstruction. We first randomly generate a set of Bézier curves to lay on the canvas, then employ the differentiable rasterizer discussed in Sec. 3.4 to render a raster image ^$\hat{\mathcal{I}}\in\mathbb{R}^{H\times W\times 3}$. The Bézier curves can then be optimized through any gradient-based loss functions. In this work, we formulate the optimization objective as minimizing the loss between $\mathcal{I}$ and $\hat{\mathcal{I}}$ while enforcing the curves to be convex. Therefore, we only adopt an $L_{2}$ loss and a Xing loss $L_{\text{Xing}}$ [17], as:
@@
Adaptive curve pruning and densification. The gradients of 2D Gaussians are influenced by local pixels only. Therefore, it is hard for 2D Gaussians to dynamically reallocate to regions requiring finer-grained details. The Gaussians would be trapped in local minima, leading to redundant Gaussians that are optimized as either low opacity or excessively large size, resulting in artifacts in rendering results.
@@
To address this issue, this work introduces a new pruning and densification strategy to dynamically adjust the density of Bézier curves throughout the optimization process. For pruning redundant Bézier curves, we apply three criteria to ensure a stable and precise optimization process. First, we remove curves with opacity below a dynamic threshold that gradually decreases as optimization progresses, ensuring that weakly contributing curves are eliminated while preserving essential structures. Second, we remove curves with an area below a predefined threshold, as they contribute minimally to the final representation. To remove visually insignificant or noisy Bézier curves, we apply an opacity-based filtering strategy: for per-segment opacities, we discard curves whose middle segment is significantly fainter than both ends—often indicating that a single curve improperly spans a region better represented by multiple curves; otherwise, we discard curves with overall opacity below a threshold (e.g., 0.2) to eliminate globally low-visibility curves. Third, we remove curves that exhibit high color similarity with surrounding curves and have significant overlap, as they provide little additional information and can be pruned without affecting the final representation. By removing such curves, we allow the optimization to introduce more appropriate replacements, leading to a more accurate and visually coherent vectorized representation. For curve densification, we adopt an error-driven curve allocation strategy inspired by LIVE [17]. Specifically, we compute connected error regions, rank them by area, and add new Bézier curves into the highest-error regions. This adaptive redistribution mechanism ensures that curves are allocated to where they would contribute the most to the reconstruction fidelity. The pruning and densification strategy enables a “global receptive field” for redistributing curve density, effectively preserving visual details for high-fidelity rendering while maintaining a compact vector representation.
Table 1: Computational speed for rendering a 2,040$\times$1,344 image with 2,048 curves.
Open curve Closed curve
DiffVG [14] Bézier Splatting Speedup DiffVG [14] Bézier Splatting Speedup
Forward pass 141.3ms 4.5ms 31.4$\times$ 85.2ms 14.1ms 6.0$\times$
Backward pass 701.3ms 4.7ms 149.2$\times$ 448.3ms 24.58ms 18.2$\times$
Figure 3: A qualitative comparison of our method and the state-of-the-art differentiable VG rasterization methods, including DiffVG [14], LIVE [17], and LIVSS [26].
@@ 4.1 Experimental Setups
Implementation details. We implement Bézier Splatting in PyTorch [19] and optimize it by using the Adam optimizer [12] with a StepLR learning rate scheduler. The learning rate is initialized at 0.01 for color, 2e-4 for Bézier curve control points, and 0.1 for opacity. For the pruning and densification strategy, the opacity threshold is set to 0.02, and the overlap threshold based on Axis-Aligned Bounding Boxes (AABB) is set to 0.9. Following the approach in LIVE [17], new curves are initialized in a circular pattern, and the number of added curves matches the number of removed ones to maintain a constant total curve count. Open curves are optimized for 15,000 iterations, and closed curves are optimized for 10,000 iterations. Pruning and densification are applied every 400 steps until the last but 1,000 steps, after which they are halted to stabilize the representation.
Datasets. We comprehensively evaluate our method across different image domains. We use the publicly available DIV2K [23] dataset for evaluating natural images. Due to the high computational cost of baseline method LIVE [17], we uniformly subsample the DIV2K dataset by selecting one out of every four images, resulting in a final evaluation set of 200 images from the original 800-image dataset. Additionally, we test our method on non-photorealistic images, including the artwork images from Clipart1K dataset [9] and cartoon images from Danbooregions dataset [31], to demonstrate its effectiveness for diverse types of images, as shown in Fig. 4.
Table 2: Quantitative evaluation and optimization efficiency of differentiable VG methods on the DIV2K dataset [23].
Method 256 curves 512 curves 1024 curves
[…]
Open DiffVG [14] 0.552 19.83 0.563 18.9min 0.587 21.47 0.537 22.0min 0.616 22.62 0.517 30.6min
Ours 0.600 22.17 0.540 3.4min 0.646 23.79 0.498 3.3min 0.699 25.45 0.448 3.2min
Closed DiffVG [14] 0.578 20.69 0.548 16.4min 0.601 21.82 0.531 18.5min 0.631 22.95 0.509 25.1min
LIVE [17] 0.576 20.09 0.543 2.6h 0.611 21.70 0.521 4.2h 0.648 23.11 0.495 5.1h
LIVSS [26] 0.586 17.71 0.542 39.2min 0.630 18.71 0.530 54.3min 0.678 19.83 0.517 1.4h
Ours 0.580 20.74 0.546 7.8min 0.607 22.11 0.528 8.3min 0.639 23.45 0.507 8.6min
Figure 4: Our Bézier Splatting achieves high-quality image vectorization results for various types of images including artworks, cartoons, and natural images. Curve type and count are indicated at the bottom right of each sample.